Let BCB′C′ be a rectangle, let M be the midpoint of B′C′, and let A be a point on the circumcircle of the rectangle. Let triangle ABC have orthocenter H, and let T be the foot of the perpendicular from H to line AM. Suppose that AM= 2, [ABC] = 2020, and BC= 10. Then AT=m/n, where m and n are positive integers with gcd (m,n) = 1. Compute 100m+n.


1
Expert's answer
2021-02-23T09:32:00-0500

"AM\cdot MK=C'M\cdot MB'\implies MK=C'M\cdot MB'/AM"


From properties of orthocenter:

"BM'=CM', HM'=A'M'=AM" , "AA'" is the diameter

Then:

"A'K'=AK,AT=HK'"


"AT=AM+MK-2AM=\frac{BC^2}{4AM}-AM"


"AT=\frac{10^2}{4\cdot2}-2=10.5=\frac{21}{2}"

"m=21,n=2"

"100m+n=100\cdot21+2=2102"